# AIDOVECL: AI-generated Dataset of Outpainted Vehicles for Eye-level Classification and Localization

### v2

@@ 3.2 Synthetic Data Generation
Seed images are randomly scaled and positioned on a $512\times 512$-pixel blank canvas. The channels of these seed images are also permuted to diversify the colorization of vehicles. However, this does not alter black and white colors, which is partly advantageous as it maintains the black color of tires. Corresponding annotations are calculated based on the dimensions of the original bounding box (i.e., dividing the buffered seed dimensions by 1.15 to remove the 15% buffer), and the scale and position of the seed on the canvas. We also create a mask image to aid the outpainting modules. This mask maintains the buffer zone—unlike the annotation which accounts for the actual bounding box of the vehicle—and will be blurred on borders to ensure a smooth transition into the outpainted areas.
@@
For the outpainting process, we employ the stable diffusion inpainting model from Hugging Face (stable_diffusion_inpainting) which is based on LDM (rombach2022high). The model utilizes both positive and negative prompts. The positive prompt is constructed dynamically according to the vehicle class and follows the pattern: "A {location} during {time} with no vehicle." The {location} parameter can take values from { highway, road, street, downtown, driveway}, tailored to the vehicle type. The {time} variable encompasses { a sunny day, a cloudy day, evening, sunset, sunrise, dusk}, providing a temporal context to complement the location. This structured prompt design ensures a contextually appropriate generation of the scene, enhancing the realism of the outpainted areas. To avoid introducing unwanted vehicles, which are not annotated in the previous step, the negative prompt explicitly excludes {traffic, train, car, truck, bus, van}. This selection relies on the fact that the masked area remains largely intact during outpainting, allowing us to maintain focus on our vehicle of the desired class. Additionally, we may add {billboard, text} to the negative prompt to avoid generating images with billboards including brands which are on the vehicles and to minimize generating advertisement-like images of vehicles. Additional terms may be added to the negative prompt to suppress artifacts that compromise the realism and quality of the generated images.
@@
We employ outpainting techniques to extend the boundaries of original vehicle images, creating visually plausible extended scenes. The quality of these outpainted images is rigorously evaluated using three no-reference image quality assessment scores: BRISQUE (mittal2012no), CLIP-IQA (wang2023exploring), and QualiCLIP (agnolucci2024quality). BRISQUE is designed to assess natural scene statistics and quantify potential distortions, typically ranging from 0 to 100, with lower scores indicating better quality. Similarly, CLIP-IQA leverages the CLIP model’s ability to assess perceptual quality through deep learning, with scores between 0 and 1 where higher scores suggest superior perceptual quality. QualiCLIP is a self-supervised, opinion-unaware image quality assessment method that fine-tunes the CLIP model to align image representations with quality-related semantics through textual prompts, outputting a score between 0 and 1 where higher values indicate better perceptual quality. We retained images with a maximum BRISQUE score of 20, a minimum CLIP-IQA score of 0.5, and a minimum QualiCLIP score of 0.85, the latter serving as a high semantic quality bar to ensure that only images with strong alignment to high-quality visual semantics were included.
@@
To generate background images for our dataset, we utilize the text-to-image stable diffusion model from Hugging Face (stable_diffusion_v1_5) based on LDM (rombach2022high). This process involves selecting from predefined descriptions that vividly depict various urban scenes devoid of any vehicles, such as empty streets and parks at different times and under various atmospheric conditions. These textual prompts guide the generation of detailed and realistic background images while using negative prompts to ensure no vehicles appear. Background images are crucial for improving the performance of visual recognition algorithms by reducing false positives and enhancing the detection of true negatives. By providing clear and unobstructed scenes, the algorithms can more accurately learn what constitutes the absence of target objects, thereby improving overall detection reliability in complex urban environments.
@@ 3.3 Vehicle Classification and Localization
In this study, we utilize the YOLOv8 (You Only Look Once) object detection framework (redmon2016you), specifically employing the YOLOv8 model (ultralytics_docs) along with four other previously mentioned models (FCOS, RetinaNet, SSD, and FasterRCNN). We begin by loading pretrained models, which provide a robust starting point due to their weights being trained on a comprehensive dataset. The training process is conducted over 1000 epochs with an early stopping patience of 20 to prevent overfitting to the training data, using a batch size of 4. The learning rates were fixed at $2\times 10^{-3}$ for FCOS, $4\times 10^{-3}$ for RetinaNet, $1\times 10^{-3}$ for SSD, and $8\times 10^{-3}$ for Faster R-CNN, while YOLOv8 used Ultralytics’ default learning-rate configuration—including a starting rate of $0.01$, final fraction of $0.01$, warm-up over 3 epochs with momentum ramping from $0.8$ and bias learning rate of $0.1$, SGD momentum of $0.937$, weight decay of $0.0005$, and a linear learning rate decay schedule with cosine one-cycle scheduling disabled by default.
@@ 4.2 Outpainted Vehicles
Samples of outpainted vehicles, as generated using the methodology described in the previous section, are illustrated in Figure 3. The diverse set of vehicle images results from varying the prompts for location and time of the scene, alongside the inherent randomness in generative inpainting models, and the seed’s scaling and positioning on the canvas. Additionally, employing no-reference image quality assessments, such as QualiCLIP, BRISQUE, and CLIP-IQA scores, ensures a degree of realism and natural appearance within the generated images. During attempts per every seed image, the method finds an outpainted image whose context is more relevant to the type and lighting of the vehicle by meeting the quality criteria.
@@
[…]
@@ 4.3 Overall Detection Performance Across Models
We conducted an ablation study that spans five modern object‑detection architectures (FCOS, RetinaNet, SSD‑Lite, FasterRCNN, and YOLOv8). For every model we evaluate four train/test configurations—Real/Real, Aug./Real, Real/Aug., and Aug./Aug.—to isolate the benefit of introducing synthetic data during training and the robustness to distribution shifts at inference. Here, augmentation signifies that each real image is accompanied with a synthetic counterpart from AIDOVECL; the generation pipeline surpasses a 91% success rate against our quality criteria after up to 100 iterations per image. Within every configuration we further vary on‑the‑fly augmentations—namely the mixup ratio and mosaic probability—to probe how aggressive blend‑and‑compose policies interact with the underlying data composition. Six standard metrics (Precision, Recall, F1, mAP50, mAP50‑95, and Fitness) are reported for all combinations.
@@
Part of the results in Tables 3, 4, and 5 compare a baseline scenario—where each detector is both trained and evaluated on real images—to a scenario where the training dataset is augmented with AIDOVECL, while evaluation remains solely on real images. That is, we first focus on "Real/Real" and "Aug./Real" results. Table 3 demonstrates that adding synthetic AIDOVECL images improves almost all metrics when no additional on-the-fly augmentations are applied. Improvements of up to 10% are noted across each model. Increasing the blending factor using Mixup (mixup=0.5) as shown in Table 4, however, does not lead to consistent or significant improvements when training data is augmented. This suggests that pixel-level blending from Mixup effectively reduces overfitting. Consequently, if inference data closely matches training distribution, substantial benefits from AIDOVECL may be limited when using Mixup simultaneously. On the other hand, Table 5 shows that when only Mosaic augmentation is applied with a probability of 0.5, there is an improvement of up to 10%. This indicates that Mosaic augmentation alone does not diminish the advantages offered by our augmentation. Therefore, when inference remains within the distribution, AIDOVECL augmentation provides clear performance benefits for most cases.
@@
Comparing the results for "Aug./Aug." and "Real/Aug." highlights the importance of augmenting the training data with AIDOVECL. Across all models, configurations, and metrics presented in Tables 3, 4, and 5, "Real/Aug." consistently underperforms compared to "Aug./Aug.". Although augmenting with AIDOVECL does not introduce entirely new vehicle objects, but instead rescales and relocates existing vehicles into new backgrounds, it remains crucial for the model to encounter these backgrounds during training to mitigate out-of-distribution inference issues. As demonstrated in all cases across the tables, "Aug./Aug." significantly outperforms "Real/Aug." by up to approximately 40%, indicating that augmenting training data with AIDOVECL effectively prepares the models for encountering vehicles in contexts different from those present in the real training dataset.
@@
Our comprehensive ablation study across multiple detection architectures clearly demonstrates the effectiveness and necessity of incorporating AIDOVECL-generated synthetic data during training. Synthetic augmentation notably enhances performance, with improvements of up to 10%. Importantly, synthetic augmentation substantially boosts robustness against distribution shifts at inference, achieving performance gains of up to approximately 40% when evaluation contexts differ from training. Thus, augmenting real datasets with AIDOVECL significantly strengthens model reliability and generalization in varied deployment scenarios.
@@ 6 Future Works
Active learning offers a promising direction for facilitating efficient seed pool construction in AIDOVECL by prioritizing which unlabeled vehicle crops—initially unclassified and manually assigned to custom categories—should be classified first. In a cold-start setting, diversity-first selection through core-set methods can ensure that the earliest labeled vehicle crops are representative of the broader data distribution (sener2017active). Task-agnostic latent-space selection, as in variational adversarial active learning, can further identify underrepresented or heterogeneous vehicle types before a classifier is trained (sinha2019variational). Once a small classifier exists, uncertainty-based querying using Bayesian methods can target ambiguous or confusing vehicle crops (gal2017deep), while batch acquisition strategies such as BADGE can balance informativeness and diversity when expanding seeds in groups (ash2019deep). In addition, loss prediction modules can highlight samples expected to be difficult for the classifier (yoo2019learning), ensuring annotation effort is concentrated where it yields the most improvement. Finally, self-training approaches can automatically assign labels to high-confidence samples while leaving manual effort for low-confidence or rare vehicle types (xie2020self). Together, these approaches could reduce manual labeling costs and accelerate seed pool growth, thereby strengthening the foundations of AIDOVECL.
@@
To overcome the limitations of single-pass multi-object outpainting—where disconnected regions compete for attention and global coherence is difficult to maintain—several promising strategies can be drawn from recent advances in conditional and controllable diffusion models: One approach is to guide generation with the original image (anchor view) and location-aware features (relative positional embeddings) (zhang2024continuous). Alternatively, a grounded outpainting model inspired by GLIGEN (li2023gligen) could condition on inputs defining the canvas layout to preserve object positions while generating coherent context. Structural inputs such as depth or edge maps, used in ControlNet (zhang2023adding), can guide outpainted regions around multiple objects to follow scene geometry and blend naturally into the environment. MultiDiffusion (bar2023multidiffusion) can be adapted for multi-object outpainting by coordinating generation across background and object-adjacent regions so that lighting, perspective, and style remain consistent. Composable diffusion methods (liu2022compositional) can improve multi-object handling by combining independent condition scores to reduce concept interference. Attend-and-Excite (xu2018attngan), while originally developed for text-to-image generation, can inspire methods that keep attention balanced across multiple objects, helping preserve coherence in their expanded surroundings. All these methods require fine-tuning of existing diffusion models to incorporate the additional conditioning and control mechanisms they rely on. Moreover, advancing data curation with ML to predict plausible object–background pairings can further enhance realism and ensure synthetic data better reflects real-world variability.
@@ 7 Conclusion
The AIDOVECL dataset represents a significant advancement in the field of machine learning for vehicle classification and localization. It addresses the critical issue of annotated data scarcity in desired classes and perspectives, and mitigates the problem of unlabeled background objects in public datasets. By incorporating generative AI techniques—specifically prompt-guided outpainting—AIDOVECL not only enriches the diversity of eye-level vehicle images but also effectively simulates realistic urban traffic scenarios. This approach is crucial for training algorithms capable of performing accurately in dynamic real-world environments. AIDOVECL’s automatic annotation, which potentially forgoes the need for manual annotation, offers a substantial improvement as demonstrated across various detection models, with or without on-the-fly augmentation methods. It can augment real datasets with class imbalances, thus enhancing the adaptability and performance of machine learning models. The dataset and its associated pipeline meet immediate training needs and fosters future advancements in computer vision technologies for applications ranging from autonomous driving to urban planning.
@@
The AIDOVECL resource is released both as (i) a dataset comprising curated and benchmark versions and (ii) a modular framework enabling end-users to generate their own augmented datasets. The dataset is openly hosted on Hugging Face at https://huggingface.co/datasets/amir-kazemi/aidovecl, accompanied by a dataset card documenting the structure, splits, and annotation format. The dataset is distributed under a CC-BY-4.0 license. The generation pipeline, with full source code and usage instructions, is publicly available on GitHub at https://github.com/amir-kazemi/aidovecl under the MIT license.
@@
In this research, we introduce a method for generating training datasets for vehicle detection and classification using outpainting techniques. The potential benefits of our work include enhanced accuracy in automated vehicle detection, which could improve urban planning, traffic management, and autonomous driving, contributing to safer and more efficient urban environments. While our dataset is carefully curated to focus on vehicle features that support these applications, we acknowledge that any form of data capture involving vehicles may contain elements that could raise privacy concerns. We emphasize the importance of ethical guidelines and responsible use of technologies to mitigate potential risks associated with data misuse. In particular, AIDOVECL is intended primarily for research and educational purposes, not for surveillance or applications that could infringe on individual rights (although because AIDOVECL only includes labels for vehicle class, we believe it to be of minimal utility for such applications). Users of the dataset and framework are expected to comply with relevant legal regulations and institutional review requirements, and to exercise caution when extending the pipeline to new domains. To support transparency, the dataset (CC-BY-4.0) and code (MIT) licenses are explicitly provided in their respective repositories.
@@
The authors acknowledge support from the University of Illinois NCSA Faculty Fellowship Program, the Zhejiang University and University of Illinois Dynamic Research Enterprise for Multidisciplinary Engineering Sciences (DREMES), and the University of Illinois Discovery Partners Institute (DPI). This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its NCSA (Gropp2023Delta).

### v3

@@ 3.2 Synthetic Data Generation
Seed images are randomly scaled and positioned on a $512\times 512$-pixel blank canvas. The channels of these seed images are also permuted to diversify the colorization of vehicles. However, this does not alter black and white colors, which is desirable as it maintains the black color of tires. Corresponding annotations are calculated based on the dimensions of the original bounding box (i.e., dividing the buffered seed dimensions by 1.15 to remove the 15% buffer), and the scale and position of the seed on the canvas. We also create a mask image to aid the outpainting modules. This mask maintains the buffer zone—unlike the annotation which accounts for the actual bounding box of the vehicle—and will be blurred on borders to ensure a smooth transition into the outpainted areas.
@@
For the outpainting process, we employ the stable diffusion inpainting model from Hugging Face (SDv1.5, 2025a) which is based on LDM (Rombach et al., 2022). The model utilizes both positive and negative prompts. The positive prompt is constructed dynamically according to the vehicle class and follows the pattern: "A {location} during {time} with no vehicle." The {location} parameter can take values from { highway, road, street, downtown, driveway}, tailored to the vehicle type. The {time} variable encompasses { a sunny day, a cloudy day, evening, sunset, sunrise, dusk}, providing a temporal context to complement the location. This structured prompt design ensures a contextually appropriate generation of the scene, enhancing the realism of the outpainted areas. To avoid introducing unwanted vehicles, which are not annotated in the previous step, the negative prompt explicitly excludes {traffic, train, car, truck, bus, van}. This selection relies on the fact that the masked area remains largely intact during outpainting, allowing us to maintain focus on our vehicle of the desired class. Additionally, we may add {billboard, text} to the negative prompt to avoid generating billboard-like artifacts when vehicle branding is treated as background advertising rather than as part of the vehicle itself. Additional terms may be added to the negative prompt to suppress artifacts that compromise the realism and quality of the generated images.
@@
We employ outpainting techniques to extend the boundaries of original vehicle images, creating visually plausible extended scenes. The quality of these outpainted images is rigorously evaluated using three no-reference image quality assessment scores: BRISQUE (Mittal et al., 2012), CLIP-IQA (Wang et al., 2023), and QualiCLIP (Agnolucci et al., 2024). BRISQUE is designed to assess natural scene statistics and quantify potential distortions, typically ranging from 0 to 100, with lower scores indicating better quality. Similarly, CLIP-IQA leverages the CLIP model’s ability to assess perceptual quality through deep learning, with scores between 0 and 1 where higher scores suggest superior perceptual quality. QualiCLIP is a self-supervised, opinion-unaware image quality assessment method that fine-tunes the CLIP model to align image representations with quality-related semantics through textual prompts, outputting a score between 0 and 1 where higher values indicate better perceptual quality. We retained images with a maximum BRISQUE score of 22, a minimum CLIP-IQA score of 0.5, and a minimum QualiCLIP score of 0.85, the latter serving as a high semantic quality bar to ensure that only images with strong alignment to high-quality visual semantics were included.
@@
To generate background images for our dataset, we utilize the text-to-image stable diffusion model from Hugging Face (SDv1.5, 2025b) based on LDM (Rombach et al., 2022). This process involves selecting from predefined descriptions that vividly depict various urban scenes devoid of any vehicles, such as empty streets and parks at different times and under various atmospheric conditions. These textual prompts guide the generation of detailed and realistic background images while using negative prompts to ensure no vehicles appear. Background images are crucial for improving the performance of visual recognition algorithms by reducing false positives and enhancing the detection of true negatives. By providing clear and unobstructed scenes, the algorithms can more accurately learn what constitutes the absence of target objects, thereby improving overall detection reliability in complex urban environments.
@@ 3.3 Vehicle Classification and Localization
In this study, we utilize the YOLOv8 (You Only Look Once) object detection framework (Redmon et al., 2016), specifically employing the YOLOv8 model (Ultralytics, 2024) along with four other previously mentioned models (FCOS, RetinaNet, SSD, and FasterRCNN). We begin by loading pretrained models, which provide a robust starting point due to their weights being trained on a comprehensive dataset. The training process is conducted over 1000 epochs with an early stopping patience of 20 to prevent overfitting to the training data, using a batch size of 4. The learning rates were fixed at $2\times 10^{-3}$ for FCOS, $4\times 10^{-3}$ for RetinaNet, $1\times 10^{-3}$ for SSD, and $8\times 10^{-3}$ for Faster R-CNN, while YOLOv8 used Ultralytics’ default learning-rate configuration—including a starting rate of $0.01$, final fraction of $0.01$, warm-up over 3 epochs with momentum ramping from $0.8$ and bias learning rate of $0.1$, SGD momentum of $0.937$, weight decay of $0.0005$, and a linear learning rate decay schedule with cosine one-cycle scheduling disabled by default.
@@ 4.2 Outpainted Vehicles
Samples of outpainted vehicles, as generated using the methodology described in the previous section, are illustrated in Figure 3. The diverse set of vehicle images results from varying the prompts for location and time of the scene, alongside the inherent randomness in generative inpainting models, and the seed’s scaling and positioning on the canvas. Additionally, employing no-reference image quality assessments, such as QualiCLIP, BRISQUE, and CLIP-IQA scores, ensures a degree of realism and natural appearance within the generated images. Across repeated attempts for each seed image, the method finds an outpainted image whose context is more relevant to the type and lighting of the vehicle by meeting the quality criteria.
@@
To quantify both the visual and semantic quality of the generated images, we report the distribution of two sets of qualities in Figure 4: 1) The same scores we used to vet the outpainted images—QualiCLIP, CLIP-IQA, and BRISQUE—satisfy the imposed thresholds of 0.85, 0.5, and 22, respectively. Moreover, mean QualiCLIP and CLIP-IQA scores are at least 0.90, with BRISQUE averaging around 13, all indicating high visual quality. 2) The sentence similarity of prompts for generation and AI-generated captions of outpainted images using existing pretrained models. For the latter, using the sentence transformer, we compare the sentence similarity between "A {vehicle class} on or in a {location} during {time}" as the expected caption and the captions predicted by BLIP and ViT-GPT2, yielding scores in [-1,1], where a score closer to 1 indicates high semantic similarity and a score closer to -1 indicates opposing semantics. In both cases, the average sentence similarity is approximately 0.5, indicating a significant and meaningful degree of semantic alignment between the generation prompts and the AI-generated captions. We additionally compare the sentence similarity between BLIP and ViT‑GPT2, as this reveals how consistently the semantics of a generated image are captured across different captioning models, providing an extra proxy for cross‑model semantic fidelity. The relatively high mean similarity of 0.7 indicates that the generated images contain semantically coherent content, as shown by consistent interpretations from distinct captioning models, which are less likely to hallucinate in different directions when the input is contextually grounded.
@@ 4.3 Overall Detection Performance Across Models
We conducted an ablation study that spans five modern object‑detection architectures (FCOS, RetinaNet, SSD‑Lite, FasterRCNN, and YOLOv8). For every model we evaluate four train/test configurations—Real/Real, Aug./Real, Real/Aug., and Aug./Aug.—to isolate the benefit of introducing synthetic data during training and the robustness to distribution shifts at inference. Here, augmentation signifies that each real image is accompanied by a synthetic counterpart from AIDOVECL; the generation pipeline surpasses a 91% success rate against our quality criteria after up to 100 iterations per image. Within every configuration we further vary on‑the‑fly augmentations—namely the mixup ratio and mosaic probability—to probe how aggressive blend‑and‑compose policies interact with the underlying data composition. Six standard metrics (Precision, Recall, F1, mAP50, mAP50‑95, and Fitness) are reported for all combinations.
@@
Part of the results in Tables 3, 4, and 5 compare a baseline scenario—where each detector is both trained and evaluated on real images—to a scenario where the training dataset is augmented with AIDOVECL, while evaluation remains solely on real images. That is, we first focus on "Real/Real" and "Aug./Real" results. Table 3 demonstrates that adding synthetic AIDOVECL images improves almost all metrics when no additional on-the-fly augmentations are applied. Improvements of up to 10% are noted across each model. Increasing mixup to 0.5 as shown in Table 4, however, does not lead to consistent or significant improvements when training data is augmented. This suggests that pixel-level blending from mixup effectively reduces overfitting. Consequently, if inference data closely matches training distribution, substantial benefits from AIDOVECL may be limited when using mixup simultaneously. On the other hand, Table 5 shows that when only mosaic augmentation is applied with a probability of 0.5, there is an improvement of up to about 10%. This indicates that mosaic augmentation alone does not diminish the advantages offered by our augmentation. Therefore, when inference remains within the distribution, AIDOVECL augmentation provides clear performance benefits for most cases.
@@
Comparing the results for "Aug./Aug." and "Real/Aug." highlights the importance of augmenting the training data with AIDOVECL. Across all models, configurations, and metrics presented in Tables 3, 4, and 5, "Real/Aug." consistently underperforms compared to "Aug./Aug.". Although augmenting with AIDOVECL does not introduce entirely new vehicle objects, it instead rescales and relocates existing vehicles into new backgrounds; therefore, it remains crucial for the model to encounter these backgrounds during training to mitigate out-of-distribution inference issues. As demonstrated in all cases across the tables, "Aug./Aug." significantly outperforms "Real/Aug." by up to around 40%, indicating that augmenting training data with AIDOVECL effectively prepares the models for encountering vehicles in contexts different from those present in the real training dataset.
@@
Our comprehensive ablation study across multiple detection architectures clearly demonstrates the effectiveness and necessity of incorporating AIDOVECL-generated synthetic data during training. Synthetic augmentation notably enhances performance, with improvements of up to about 10%. Importantly, synthetic augmentation substantially boosts robustness against distribution shifts at inference, achieving performance gains of up to around 40% when evaluation contexts differ from training. Thus, augmenting real datasets with AIDOVECL significantly strengthens model reliability and generalization in varied deployment scenarios.
@@ 6 Future Work
Active learning offers a promising direction for facilitating efficient seed pool construction in AIDOVECL by prioritizing which unlabeled vehicle crops—initially unclassified and manually assigned to custom categories—should be classified first. In a cold-start setting, diversity-first selection through core-set methods can ensure that the earliest labeled vehicle crops are representative of the broader data distribution (Sener and Savarese, 2017). Task-agnostic latent-space selection, as in variational adversarial active learning, can further identify underrepresented or heterogeneous vehicle types before a classifier is trained (Sinha et al., 2019). Once a small classifier exists, uncertainty-based querying using Bayesian methods can target ambiguous or confusing vehicle crops (Gal et al., 2017), while batch acquisition strategies such as BADGE can balance informativeness and diversity when expanding seeds in groups (Ash et al., 2019). In addition, loss prediction modules can highlight samples expected to be difficult for the classifier (Yoo and Kweon, 2019), ensuring annotation effort is concentrated where it yields the most improvement. Finally, self-training approaches can automatically assign labels to high-confidence samples while leaving manual effort for low-confidence or rare vehicle types (Xie et al., 2020). Together, these approaches could reduce manual labeling costs and accelerate seed pool growth, thereby strengthening the foundations of AIDOVECL.
@@
To overcome the limitations of single-pass multi-object outpainting—where disconnected regions compete for attention and global coherence is difficult to maintain—several promising strategies can be drawn from recent advances in conditional and controllable diffusion models: One approach is to guide generation with the original image (anchor view) and location-aware features (relative positional embeddings) (Zhang et al., 2024). Alternatively, a grounded outpainting model inspired by GLIGEN (Li et al., 2023) could condition on inputs defining the canvas layout to preserve object positions while generating coherent context. Structural inputs such as depth or edge maps, used in ControlNet (Zhang et al., 2023a), can guide outpainted regions around multiple objects to follow scene geometry and blend naturally into the environment. MultiDiffusion (Bar-Tal et al., 2023) can be adapted for multi-object outpainting by coordinating generation across background and object-adjacent regions so that lighting, perspective, and style remain consistent. Composable diffusion methods (Liu et al., 2022) can improve multi-object handling by combining independent condition scores to reduce concept interference. Attend-and-Excite (Xu et al., 2018), while originally developed for text-to-image generation, can inspire methods that keep attention balanced across multiple objects, helping preserve coherence in their expanded surroundings. All these methods require fine-tuning of existing diffusion models to incorporate the additional conditioning and control mechanisms they rely on. Moreover, advancing data curation with ML to predict plausible object–background pairings can further enhance realism and ensure synthetic data better reflects real-world variability.
@@ 7 Conclusion
The AIDOVECL dataset represents a significant advancement in the field of machine learning for vehicle classification and localization. It addresses the critical issue of annotated data scarcity in desired classes and perspectives, and mitigates the problem of unlabeled background objects in public datasets. By incorporating generative AI techniques—specifically prompt-guided outpainting—AIDOVECL not only enriches the diversity of eye-level vehicle images but also effectively simulates realistic urban traffic scenarios. This approach is crucial for training algorithms capable of performing accurately in dynamic real-world environments. AIDOVECL’s automatic annotation, which potentially mitigates the need for manual annotation, offers a substantial improvement as demonstrated across various detection models, with or without on-the-fly augmentation methods. It can augment real datasets with class imbalances, thus enhancing the adaptability and performance of machine learning models. The dataset and its associated pipeline meet immediate training needs and foster future advancements in computer vision technologies for applications ranging from autonomous driving to urban planning.
@@
The AIDOVECL resource is released as both (i) a dataset, comprising curated and benchmark versions, and (ii) a modular framework for generating customized augmented datasets. The dataset is publicly hosted on Hugging Face, with an accompanying dataset card describing its structure, splits, and annotation format, and is distributed under the CC-BY-4.0 license, and is archived under DOI 10.57967/hf/8444 (Kazemi et al., 2026). The associated generation pipeline, including source code and usage instructions, is publicly available on GitHub under the MIT license:
https://huggingface.co/datasets/amir-kazemi/aidovecl
https://github.com/amir-kazemi/aidovecl
@@
In this research, we introduce a method for generating training datasets for vehicle detection and classification using outpainting techniques. The potential benefits of our work include enhanced accuracy in automated vehicle detection, which could improve urban planning, traffic management, and autonomous driving, contributing to safer and more efficient urban environments. While our dataset is carefully curated to focus on vehicle features that support these applications, we acknowledge that any form of data capture involving vehicles may contain elements that could raise privacy concerns. We emphasize the importance of ethical guidelines and responsible use of technologies to mitigate potential risks associated with data misuse. In particular, AIDOVECL is intended primarily for research and educational purposes, not for surveillance or other uses that could infringe on individual rights, especially since it includes labels only at the vehicle-class level rather than finer-grained vehicle attributes, thereby limiting its utility for such purposes. Users of the dataset and framework are expected to comply with relevant legal regulations and institutional review requirements, and to exercise caution when extending the pipeline to new domains. To support transparency, the dataset (CC-BY-4.0) and code (MIT) licenses are explicitly provided in their respective repositories.
@@
The authors acknowledge support from the University of Illinois NCSA Faculty Fellowship Program, the Zhejiang University and University of Illinois Dynamic Research Enterprise for Multidisciplinary Engineering Sciences (DREMES), and the University of Illinois Discovery Partners Institute (DPI). This research used the Delta advanced computing and data resource which is supported by the National Science Foundation (award OAC 2005572) and the State of Illinois. Delta is a joint effort of the University of Illinois Urbana-Champaign and its NCSA (Gropp et al., 2023).